# Emergence of the Primacy Effect in Structured State-Space Models

### v3

@@ 2.2 Structured State-Space Models
This flexibility extends the applicability of SSMs to inherently discrete data lacking overt continuous dynamics (e.g., text languages); the model can jointly lean latent representations (or embeddings) of discrete inputs along with their (pseudo-)continuous dynamics via gradient-based optimization. Furthermore, SSMs can be hierarchically stacked to construct deeper and more expressive ANNs, where each layer processes latent signals received from lower layers.
Figure 1: Illustration of the continuous (left; $K(-t):=e^{tA}B$) and discretized (right; $\bar{K}_{-j}:=\bar{A}^{j}\bar{B}$) SSM kernels based on Legendre polynomials with an exponentially decaying measure (Gu et al.,, 2023). For a more intuitive visualization (particularly for readers unfamiliar with convolutional operations), the sign of the horizontal axis has been flipped (ranging in the negative domain) so that kernel values multiplied with past inputs appear on the left (unlike in the standard visualization of convolutional kernels, where they are placed on the right). The polynomial degree is indicated by the line colors. The kernels were evaluated at $t=j\Delta t$ for $j=0,\dots,255$ and $\Delta t\in\{0.001,0.01,0.1,1.0\}$. Increasing $\Delta t$ results in growing discrepancies between the continuous and discretized kernels.
@@ 2.3 Finite-Queue/Markov Models
Finally, it is worth mentioning finite-queue/Markov models, which encompass the current goldstandard architecture of ANNs, Transformer (Vaswani et al.,, 2017). These models are equipped with an explicit storage mechanism that retains raw observation records up to a prespecified limit. Formally, they enable arbitrary computations at the current discrete time $j$ while maintaining full access to the most recent $M$ observations, $x_{j-M+1},\dots,x_{j}$. At each time step, they “forget” or discard the oldest observation ($x_{j-M}$) as new data enter the storage buffer.
By definition, finite-queue/Markov models do not exhibit the primacy effect, in a sense that the earliest observations are inevitably lost once they exceed the finite storage window. Instead, these models align with the classical view of biological short-term memory Miller, (1956), which has traditionally been linked with the recency effect (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968). Given this fundamental limitation, finite-queue/Markov models fall outside the scope of the present study.
@@ 3.1 Task
[…]
The memorization patterns of ANNs were assessed using the binary memory verification task (Fig. 2; a.k.a. serial probe recognition; Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985).444 The most widely adopted task for assessing the primacy effect in human memory is free recall, in which participants are presented with a sequence of items and subsequently asked to recall them in an order-agnostic manner (Murdock,, 1962; Glanzer and Cunitz,, 1966). While this paradigm can be technically formulated as a loss function for ANNs (Cuturi,, 2013), initial explorations of this study revealed that model performance remained suboptimal under this approach, yielding lower accuracy than in theoretically more demanding tasks requiring order-sensitive reconstruction. Consequently, the present study adopted the more machine learning-friendly task based on binary verification. Remarkably, this task has also been used to assess the memory capacity of non-human animals, which are unable to perform free recall (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). In this task, the models were first presented with a sequence of randomly generated, non-repeating integers (hereinafter referred to as study items). Subsequently, they received another sequence of integer queries and were trained to determine whether each query token was present (labeled as 1) or absent (labeled as 0) in the study items. To construct these queries, the study items were first shuffled, and then, with a probability of $p=0.5$, each shuffled token was replaced with a randomly sampled integer from the complement set of the study items (termed distractors).
@@ 3.2 Models
This study primarily examined the single-layer S4 model as the goldstandard implementation of the SSM (Gu et al., 2022b, ).555 Recent studies have shown that the state matrix ($A$) of S4 can be simplified into a purely diagonal form without compromising performance (S4D; Gu et al., 2022a, ). By contrast, the original S4 model introduced an additional low-rank component to the diagonal structure (referred to as the Diagonal Plus Low Rank form, or DPLR) to ensure a mathematically well-founded state matrix. Notably, the diagonal variant exhibited a qualitatively similar primacy effect to the DPLR model. Due to the page limitations, results for the diagonal model are omitted from this paper, and all reported findings are based on the DPLR model. The model encoded the channel-wise dynamics of the input embeddings in a complex-valued space, with its outputs subsequently projected back into the real domain by discarding imaginary components. The state and input matrices ($A$ and $B$ in Eq. 1) were initialized to approximate each channel’s trajectory using Legendre/Laguerre polynomials of degrees 0–63 (HiPPO-LegS/LagT) or a Fourier basis $\{s_{0},c_{0},\dots,s_{31},c_{31}\}$, where $s_{n}(t):=\sqrt{2}\sin(2\pi nt)$ and $c_{n}(t):=\sqrt{2}\cos(2\pi nt)$ (HiPPO-Fout, Fourier Recurrent Unit; Zhang et al.,, 2018; Gu et al.,, 2020, 2023). The matrices were discretized by the bilinear method (Tustin,, 1947).
@@
Figure 3: Accuracy of the binary memory verification task. Each cell in the heatmap represents the accuracy for study items that were first observed at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. The accuracy for distractor queries is displayed in the top separate row of each panel, alongside the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across verification times (columns). The bottom-right panel (3) depicts the accuracy distribution for the LSTM, while the other panels (33) report results for the S4 model under different parameter configurations. The state and input matrices of the SSM were initialized to approximate the latent dynamics of input sequences using Legendre polynomials, except in panels 3 and 3, where Laguerre and Fourier bases were used, respectively. The state and input matrices were optimized for the task in panel 3, whereas they remained fixed at their initial values in all other panels. The discretization step size $\Delta t$ was initialized in the range $0.001\leq\Delta t\leq 0.1$, except in panel 3, where the upper bound was extended to $1.0$ (i.e., $0.001\leq\Delta t\leq 1.0$). The number of study items was set to $L=128$, except in panel 3 ($L=64$) and panel 3 ($L=256$).
Figure 4: Top row: Optimization trajectories of the discretization step size $\Delta t$ in the S4 model with frozen Legendre state and input matrices. Each heatmap column represents the distribution of $\Delta t$ across 256 latent channels $\times$ 10 training runs, sorted in the ascending order. Bottom row: Histograms displaying the initial (blue) and final (orange) values of $\Delta t$. The first to third panel columns from the left show the results for three different ranges of log-uniformly random initializations, whereas the fourth and fifth columns tested shorter and longer study items, respectively.
4 Results
4.1 Emergence of the Primacy Effect
The binary memory verification performance of the S4 model was highest for study items presented at the beginning of the sequence, demonstrating a clear primacy effect (Fig. 33). Additionally, the model maintained high accuracy across different query timings (as indicated by the bright colors in the top rows of the heatmaps), provided that the sequence length did not exceed its capacity (see the accuracy decline in Fig. 3, where $L=256$). In other words, memory for the initial study items exhibited minimal decay over time.
Interestingly, the S4’s accuracy for the most recently presented study items was lowest when they were queried immediately after their initial presentation in the memorization phase (indicated by the dark colors in the bottom-left region of the heatmaps). This suggests a temporal delay between the encoding of study items and their effective retrieval.
@@ 4.2 Distribution of the Time-Step Sizes
As discussed in the Preliminaries, the discretization time-step size $\Delta t$ is a key parameter governing the memory capacity of the SSM, although the primacy effect emerged across different initial values (Fig. 3). To further investigate its role, the learning trajectories of $\Delta t$ over training iterations were analyzed. The results indicated that as training progressed, a specific range of step sizes ($\Delta t\leq 0.03$) became dominant, compensating for the deallocation of the higher range approximately between 0.03 and 0.2 (Fig 4).
@@ 5.1 Summary of the Findings
However, little attention has been given to how the model handles longer study sequences and larger vocabularies when memory capacity reaches its limit.666 It should be noted that the performance of the SSM can be enhanced by increasing the number of layers and/or latent channels. In this study, the model’s capacity was intentionally constrained in order to study its behavior under conditions where perfect accuracy is unattainable. In particular, the question of whether the models prioritize initial/middle/recent observations has remained unexplored. The present study addressed this question and discovered that the SSM predominantly preserved the initial observations.
The key factor responsible for the primacy effect in the SSM appears to be the time-step size, $\Delta t$. After training on the memorization task, $\Delta t$ values concentrated below a specific threshold ($\Delta t\leq 0.03$). As discussed in the Preliminaries, smaller $\Delta t$ values allow the model to retain more distant memories, while larger $\Delta t$ values enhance the discrimination of adjacent tokens (Gu et al.,, 2021; Gu and Dao,, 2024). The learning results thus align with the necessary condition for the primacy effect; however, the question remains as to why recent observations were remembered less accurately despite the exponentially decaying measure underlying the polynomial-approximation theory.
5.2 Comparison with Data/Task-Oriented Primacy Effects
Several prior studies have reported a primacy effect of ANNs trained on the autoregressive language modeling task. Wang et al., (2023) examined positional biases in a Transformer-based large language model (LLM) using a prompting-based approach. Specifically, a list of action or event labels was sequentially presented (e.g., “Label 1: change_pin”, “Label 2: card_arrival”, “Label 3: activate_my_card”). The model was then given a query prompt specifying a target action/event (e.g., “Target Text: I need a new PIN.”) along with an instruction statement (e.g., “Which label matches the intent expressed in the Target Text?”). The primacy effect was observed as a greater frequency of the initially presented labels in the model’s responses.777 According to the published code, the indices assigned to the labels (e.g., “2” in “Label 2: card_arrival”) consistently corresponded to their order of presentation; that is, “Label $j$: xxx” was never listed before “Label $j^{\prime}$: yyy” if $j^{\prime}<j$, even though the label contents (“xxx” and “yyy”) were disassociated from the label indices and ordered randomly. Hence, the reported primacy effect could be attributed to a bias in model responses toward lower-indexed labels, which appear more frequently than larger integers (e.g., 2 vs. 67), rather than to greater attentions to initially presented labels.
Xiao et al., (2024) found that Transformer-based LLMs allocated disproportionate attention to initial tokens, regardless of their informational salience in the text (a phenomenon they termed attention sinks). Furthermore, retaining these initial tokens even after they fall outside the prespecified input window enhanced model performance.
By contrast, the findings of the present study are based on a synthetic memorization task designed to closely resemble controlled psychological experiments (Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). Therefore, the observed primacy effect is attributed to the intrinsic properties of the SSM per se, rather than to biases introduced by data or task design.
@@ 5.3 Implications for Computational Neuroscience
Classical theories on the primacy effect in human and animal memory hypothesize a dual-storage mechanism, attributing the primacy and recency effects to long- and short-term memory systems, respectively (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968). According to these theories, initially presented study items are transferred from short-term memory—where all incoming observations are temporally buffered—into long-term memory, where they are preserved against decay, thereby explaining for the primacy effect. By contrast, the recency effect arises from the residual information retained in short-term memory, which has a limited capacity and can store the most recently presented items.
By contrast, the findings of the present study may contribute to the development of a novel and more neurobiologically grounded model of the primacy effect. Voelker and Eliasmith, (2018) and Voelker et al., (2019) established a connection between the SSM and spiking neural networks, demonstrating that the dynamics of the state vector in the SSM (Eq. 1) can be implemented using a population of spiking neurons. Given that spiking neural networks remain the most plausible model of biological neuronal activity to date, a spiking version of the SSM studied here could serve as a new neurobiological model of the primacy effect.
Furthermore, Voelker and Eliasmith, (2018) also showed that the spiking representation of an SSM, when tuned to implement some particular function known as the delay line, replicates the behavior of biological time cells,888 Time cells fire at specific moments following onset cues/events and are believed to encode temporal information (MacDonald et al.,, 2011). which have been observed in several brain regions, including the hippocampus (MacDonald et al.,, 2011; Eichenbaum,, 2014), striatum (Mello et al.,, 2015), and cortex (Luczak et al.,, 2015). In such a way, comparing a spiking implementation of the SSM that exhibits the primacy effect with brain recordings from psychological experiments could provide novel computational insights into the neuronal mechanisms underlying this cognitive bias.
@@
The model’s capacity also far exceeded that of biological memory systems. The binary memory verification task employed in this study involved extremely long lists of study items ($L\in\{64,128,256\}$) as well as a huge vocabulary ($K=4096$), which is unlikely to be replicated in human or animal experiments. This level of task difficulty was necessary to clearly demonstrate the primacy effect in the SSM. However, future research should explore downscaling the model toward more biologically plausible conditions.

### v4

@@ 2.2 Structured State-Space Models
This flexibility extends the applicability of SSMs to inherently discrete data lacking overt continuous dynamics (e.g., text languages; Gu and Dao,, 2024; Dao and Gu,, 2024); the model can jointly lean latent representations (or embeddings) of discrete inputs along with their (pseudo-)continuous dynamics via gradient-based optimization. Furthermore, SSMs can be hierarchically stacked to construct deeper and more expressive ANNs, where each layer processes latent signals received from lower layers.
Figure 1: Illustration of the continuous (left; $K(\tau):=e^{\tau A}B$) and discretized (right; $\bar{K}_{j}:=\bar{A}^{j}\bar{B}$) SSM kernels based on Legendre polynomials with an exponentially decaying measure (Gu et al.,, 2023). To aid intuitive understanding (particularly for readers unfamiliar with convolutional operations), the horizontal axis has been flipped, so that kernel values multiplied with past inputs appear on the left (unlike in the standard visualization of convolutional kernels, where they are placed on the right). The distinct line colors represent selected entries of the kernel vectors, $K_{N}(\tau)$ and $\bar{K}_{j,n}$, each corresponding the approximating polynomial of degree $n\in\{2,3,8,15,32,63\}$. The kernels were evaluated at $\tau=j\Delta t$ for $j=0,\dots,255$ and $\Delta t\in\{0.001,0.01,0.1,1.0\}$. Increasing $\Delta t$ results in growing discrepancies between the continuous and discretized kernels.
@@ 2.3 Finite-Queue/Markov Models
Finally, it is worth mentioning finite-queue/Markov models, which encompass the current goldstandard architecture of ANNs, Transformer (Vaswani et al.,, 2017). These models are equipped with an explicit storage mechanism that retains raw observation records up to a prespecified limit. Formally, they enable arbitrary computations at a given discrete time $j$ while maintaining full access to the most recent $M$ observations, $x_{j-M+1},\dots,x_{j}$. At each time step, the oldest observation ($x_{j-M}$) is discarded as new data enter the storage buffer.
By design, finite-queue/Markov models do not exhibit the primacy effect once earlier observations inputs fall outside the accessible memory window. Instead, they align with the classical view of biological short-term memory Miller, (1956), which has traditionally been linked with the recency effect (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968).
Nonetheless, given sufficient computational resources, the memory capacity of modern models like Transformers can cover most empirical sequences, rendering the out-of-window effects negligible in practice. In such cases, models with naive implementations assign equal importance to all observations. Most notably, the canonical Transformer is inherently agnostic to input order in the absence of positional encoding (Vaswani et al.,, 2017), implying a form of bias-free memorization. Nevertheless, as discussed in the following section, even such unbiased models can exhibit the primacy effect when trained on specific datasets or tasks.
2.4 Data-Driven Primacy Effect in Language Models
Several prior studies have documented the primacy effect of ANNs trained on the autoregressive language modeling task. Wang et al., (2023) investigated positional biases in a Transformer-based large language model (LLM) using a prompting-based approach. Specifically, a list of action or event labels was sequentially presented (e.g., “Label 1: change_pin”, “Label 2: card_arrival”, “Label 3: activate_my_card”). The model was then given a query prompt specifying a target action/event (e.g., “Target Text: I need a new PIN.”) along with an instruction statement (e.g., “Which label matches the intent expressed in the Target Text?”). The primacy effect was observed as a greater frequency of the initially presented labels in the model’s responses.444 According to the published code, the indices assigned to the labels (e.g., “2” in “Label 2: card_arrival”) consistently corresponded to their order of presentation; that is, “Label $j$: xxx” was never listed before “Label $j^{\prime}$: yyy” if $j^{\prime}<j$, even though the label contents (“xxx” and “yyy”) were disassociated from the label indices and ordered randomly. Hence, the reported primacy effect could be attributed to a bias in model responses toward lower-indexed labels, which appear more frequently than larger integers (e.g., 2 vs. 67), rather than to greater attentions to initially presented labels. Comparable findings have been reported across different LLM implementations and benchmark datasets (Eicher and Irgolič,, 2024; Guo and Vosoughi,, 2024; Janik,, 2024; Liu et al.,, 2024).
Xiao et al., (2024) found that Transformer-based LLMs allocated disproportionate attention to initial tokens, regardless of their informational salience in the text (a phenomenon they termed attention sinks). Furthermore, retaining these initial tokens even after they fall outside the predefined input window was found to enhance model performance.
[…]
In contrast to prior investigations, the present study examines the emergence of the primacy effect in ANNs preventing the inheritance of human-induced bias. Specifically, the models are trained on a synthetic memorization task designed based on psychological experiments conducted with humans and other animals (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). The following section details the task formulation and the model architecture.
@@ 3.1 Task
The memorization patterns of ANNs were assessed using the binary memory verification task (Fig. 2; a.k.a. serial probe recognition; Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985).555 The most widely adopted task for assessing the primacy effect in human memory is free recall, in which participants are presented with a sequence of items and subsequently asked to recall them in an order-agnostic manner (Murdock,, 1962; Glanzer and Cunitz,, 1966). While this paradigm can be technically formulated as a loss function for ANNs (Cuturi,, 2013), initial explorations of this study revealed that model performance remained suboptimal under this approach, yielding lower accuracy than in theoretically more demanding tasks requiring order-sensitive reconstruction. Consequently, the present study adopted the more machine learning-friendly task based on binary verification. Remarkably, this task has also been used to assess the memory capacity of non-human animals, which are unable to perform free recall (Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). In this task, the models were first presented with a sequence of randomly generated, non-repeating integers (hereinafter referred to as study items). Subsequently, they received another sequence of integer queries and were trained to determine whether each query token was present (labeled as 1) or absent (labeled as 0) in the study items. To construct these queries, the study items were first shuffled, and then, with a probability of $p=0.5$, each shuffled token was replaced with a randomly sampled integer from the complement set of the study items (termed distractors).
@@ 3.2 Models
This study primarily examined the single-layer S4 model as the goldstandard implementation of the SSM (Gu et al., 2022b, ).666 Recent studies have shown that the state matrix ($A$) of S4 can be simplified into a purely diagonal form without compromising performance (S4D; Gu et al., 2022a, ). By contrast, the original S4 model introduced an additional low-rank component to the diagonal structure (referred to as the Diagonal Plus Low Rank form, or DPLR) to ensure a mathematically well-founded state matrix. Notably, the diagonal variant exhibited a qualitatively similar primacy effect to the DPLR model. Due to the page limitations, results for the diagonal model are omitted from this paper, and all reported findings are based on the DPLR model. The model encoded the channel-wise dynamics of the input embeddings in a complex-valued space, with its outputs subsequently projected back into the real domain by discarding imaginary components. The state and input matrices ($A$ and $B$ in Eq. 1) were initialized to approximate each channel’s trajectory using Legendre/Laguerre polynomials of degrees 0–63 (HiPPO-LegS/LagT) or a Fourier basis $\{s_{0},c_{0},\dots,s_{31},c_{31}\}$, where $s_{n}(t):=\sqrt{2}\sin(2\pi nt)$ and $c_{n}(t):=\sqrt{2}\cos(2\pi nt)$ (HiPPO-Fout, Fourier Recurrent Unit; Zhang et al.,, 2018; Gu et al.,, 2020, 2023). The matrices were discretized by the bilinear method (Tustin,, 1947).
@@
Another well-established baseline worth considering is the Transformer architecture (Vaswani et al.,, 2017), equipped with a sufficiently long memory window. However, it turned out that the Transformer can easily achieve perfect accuracy on the adopted task—even at the maximal levels of input length and vocabulary size implementable within the available computational resources—thereby failing to incur the memory load necessary for evaluating the primacy effect777 For the same reason, the proposed experimental paradigm was also inadequate for testing extended architectures of the SSM (Mamba; Gu and Dao,, 2024; Dao and Gu,, 2024) and the LSTM (xLSTM; Beck et al.,, 2024), both of which attained perfect accuracy at the highest task difficulty settings. Consequently, the Transformer is excluded from the main results reported in the next section; instead, the Appendix provides a detailed explanation of how it can solve the memory verification task, with particular emphasis on the role of its attention mechanism.
@@
4 Results
4.1 Emergence of the Primacy Effect
Fig. 3 reports the accuracy of the binary memory verification task across all combinations of memorization and verification times. The brightness of each cell in the square heatmaps indicates the accuracy for study items that were presented at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. That is, they report the proportion of true positives against false negatives (i.e., the recall score). Additionally, the top separate row of each panel displays the accuracy for distractor queries (integers not included among the study items), capturing the prevalence of true negatives over false positives. Just below it, the second row summarizes the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across all verification times (columns).
The binary memory verification performance of the SSM model was highest for study items presented at the beginning of the sequence, demonstrating a clear primacy effect (Fig. 33). The model maintained high accuracy across different query timings (as indicated by the bright colors in the top rows of the heatmaps), provided that the sequence length did not exceed its capacity (see the accuracy decline in Fig. 3, where $L=256$). In other words, memory for the initial study items exhibited minimal decay over time.
Interestingly, the SSM’s accuracy for the most recently presented study items was lowest when they were queried immediately after their initial presentation in the memorization phase (indicated by the dark colors in the bottom-left region of the heatmaps). This suggests a temporal delay between the encoding of study items and their effective retrieval.
@@ 4.1 Emergence of the Primacy Effect
Figure 3: Accuracy of the binary memory verification task. Each cell in the square heatmaps represents the accuracy (or the recall score) for study items that were presented at the time indexed by the corresponding row and queried at the time indexed by the corresponding column. The accuracy for distractor queries is displayed in the top separate row of each panel, alongside the average accuracy across memorization times (rows). Similarly, the rightmost separate column represents the average accuracy across verification times (columns). The bottom-right panel (3) depicts the accuracy distribution for the LSTM, while the other panels (33) report results for the SSM (S4) model under different parameter configurations. The state and input matrices of the SSM were initialized to approximate the latent dynamics of input sequences using Legendre polynomials, except in panels 3 and 3, where Laguerre and Fourier bases were used, respectively. The state and input matrices were optimized for the task in panel 3, whereas they remained fixed at their initial values in all other panels. The discretization step size $\Delta t$ was initialized in the range $0.001\leq\Delta t\leq 0.1$, except in panel 3, where the upper bound was extended to $1.0$ (i.e., $0.001\leq\Delta t\leq 1.0$). The number of study items was set to $L=128$, except in panel 3 ($L=64$) and panel 3 ($L=256$).
Figure 4: Top row: Optimization trajectories of the discretization step size $\Delta t$ in the SSM (S4) model with frozen Legendre state and input matrices. Each heatmap column represents the distribution of $\Delta t$ across 256 latent channels $\times$ 10 training runs, sorted in the ascending order. Bottom row: Histograms displaying the initial (blue) and final (orange) values of $\Delta t$, aggregated across the 256 latent channels $\times$ 10 training runs (see Fig. 5 in the Appendix for variations among individual runs). The first to third panel columns from the left show the results for three different ranges of log-uniformly random initializations, whereas the fourth and fifth columns tested shorter and longer study items, respectively.
@@ 4.2 Distribution of the Time-Step Sizes
As discussed in the Preliminaries, the discretization time-step size $\Delta t$ plays a critical role in determining the memory capacity of the SSM. Moreover, once the state matrix $A$ is fixed, $\Delta t$ becomes the sole parameter capable of influencing the dynamics of the SSM;888 Freezing $\Delta t$ resulted in a complete failure of learning. all remaining parameters are confined to feedforward transforms.
Accordingly, to further investigate its role, the optimization trajectories of $\Delta t$ were tracked over the course of training. The analysis revealed that as training progressed, a specific range of step sizes ($\Delta t\leq 0.03$) became dominant, compensating for the deallocation of the higher range approximately between 0.03 and 0.2 (Fig. 4).
@@ 5.1 Summary of the Findings
However, little attention has been given to how the model handles longer study sequences and larger vocabularies when memory capacity reaches its limit.999 It should be noted that the performance of the SSM can be enhanced by increasing the number of layers and/or latent channels. In this study, the model’s capacity was intentionally constrained in order to study its behavior under conditions where perfect accuracy is unattainable. In particular, the question of whether the models prioritize initial/middle/recent observations has remained unexplored. The present study addressed this question and discovered that the SSM predominantly preserved the initial observations.
The key factor responsible for the primacy effect in the SSM appears to be the time-step size, $\Delta t$, as all the other trainable parameters pertain exclusively to feedforward transforms. After training on the memorization task, $\Delta t$ values concentrated below a specific threshold ($\Delta t\leq 0.03$). As discussed in the Preliminaries, smaller $\Delta t$ values allow the model to retain more distant memories, while larger $\Delta t$ values enhance the discrimination of adjacent tokens (Gu et al.,, 2021; Gu and Dao,, 2024). The learning results thus align with the necessary condition for the primacy effect; however, the question remains as to why recent observations were remembered less accurately despite the exponentially decaying measure underlying the polynomial-approximation theory.
The SSMs analyzed in this study were trained from scratch on a synthetic memorization task that was designed to closely resemble controlled psychological experiments (Wickelgren and Norman,, 1966; Thompson and Herman,, 1977; Sands and Wright,, 1980; Wright et al.,, 1985). Consequently, the observed primacy effect is attributed to the intrinsic properties of the SSM per se, rather than to biases introduced by data or task design. From this perspective, the the present study stands in contrast to prior investigations of LLMs (Wang et al.,, 2023; Eicher and Irgolič,, 2024; Guo and Vosoughi,, 2024; Janik,, 2024; Liu et al.,, 2024; Xiao et al.,, 2024); LLMs are trained on human-generated linguistic data and therefore likely to inherit the primacy effect as a byproduct of human cognitive biases embedded in the data.
@@ 5.2 Implications for Computational Neuroscience
Most existing theories of the primacy effect in human and animal memory have focused on abstract, procedural explanations of the cognitive bias. One influential account is the dual-store hypothesis (Waugh and Norman,, 1965; Glanzer and Cunitz,, 1966; Atkinson and Shiffrin,, 1968), which attributes the primacy and recency effects to two distinct memory systems. According to this view,study items presented early in a sequence are transferred from short-term memory—where all incoming observations are temporally buffered—into long-term memory, where they are preserved against decay, thereby producing the primacy effect. By contrast, the recency effect is thought to arise from the residual information retained in short-term memory, which has limited capacity and only holds the most recent inputs.
Moving beyond such abstract accounts, several researchers have proposed biologically grounded neural network models that replicate the primacy effect (Burgess et al.,, 1991; Wong et al.,, 1991; Greene et al.,, 2000; Sikström,, 2006; Lansner et al.,, 2013). The present findings align with this line of work but offer a novel perspective. Specifically, the SSM examined here was developed with the aim of advancing practically useful artificial intelligence, independently of any deliberate attempt to model the primacy effect from a psychological/biological standpoint. Thus, the spontaneous emergence of the primacy effect in the SSM presents a unique and complementary hypothesis for the mechanism underlying this cognitive bias, opening new avenues for future investigation in computational neuroscience.
Importantly, the industrial orientation of the SSM does not imply biological implausibility. Voelker and Eliasmith, (2018) and Voelker et al., (2019) established a connection between the SSM and spiking neural networks, demonstrating that the dynamics of the state vector in the SSM (Eq. 1) can be implemented using a population of spiking neurons, which remain the most plausible model of biological neuronal activity to date. Moreover, Voelker and Eliasmith, (2018) showed that the spiking instantiation of the SSM, when tuned to implement some particular function known as the delay line, reproduces the characteristic behavior of biological time cells,101010 Time cells fire at specific moments following onset cues/events and are believed to encode temporal information (MacDonald et al.,, 2011). which have been observed in several brain regions, including the hippocampus (MacDonald et al.,, 2011; Eichenbaum,, 2014), striatum (Mello et al.,, 2015), and cortex (Luczak et al.,, 2015). Comparing a spiking implementation of the SSM exhibiting the primacy effect with brain recordings from psychological experiments may thus yield novel insights into the neuro-computational mechanisms underlying this cognitive phenomenon.
@@
It is also important to note that the SSM does not constitute a comprehensive model of sequence memorization; its lack of the recency effect suggests the necessity of incorporating an additional module to account for short-term memory processes. Accordingly, future work should investigate whether the SSM can reliably function as a long-term memory component when integrated with a complementary short-term memory mechanism (e.g., finite-queue/Markov systems), either through modular combination or ensemble learning.
The model’s capacity also far exceeded that of biological memory systems. The binary memory verification task employed in this study involved extremely long lists of study items ($L\in\{64,128,256\}$) as well as a huge vocabulary ($K=4096$), which is unlikely to be replicated in human or animal experiments. This level of task difficulty was necessary to clearly demonstrate the primacy effect in the SSM, and was still insufficient to reach the memory limits of more powerful architectures such as Mamba (Gu and Dao,, 2024; Dao and Gu,, 2024) and xLSTM (Beck et al.,, 2024),thereby constraining the scope of model comparison here. Future research should explore downscaling the model toward more biologically plausible conditions.